# Continuous time asymptotic representations for adaptive experiments

### v1

@@ A.2. Proof of Theorem 1
We can informally outline the proof as follows: We employ dyadic approximations to discretize the empirical allocation processes $\{q_{n,a}(\cdot)\}_{a}$ across both their support (i.e., time) and range. The resulting discrete processes are shown to converge in distribution under $\mathbb{P}_{n,0}$ to a limit (discrete) process that is a function of $(z_{1}(\cdot),z_{0}(\cdot),U)$, where $z_{1}(\cdot),z_{0}(\cdot)$ are independent $d$-dimensional Wiener processes, and $U\sim\textrm{Uniform}[0,1]$ represents exogenous randomization.
Next, we demonstrate that as the discretization becomes arbitrarily fine, the intermediate limit processes converge to a continuous-time allocation process satisfying the criteria in Definition 1. The theory of stable convergence (hausler2015stable) plays a key role here in ensuring that the informational/measurability constraints satisfied by the empirical allocation processes are preserved during the transition to the continuous-time (and range) limit.
@@
Step 1 (Convergence under dyadic approximations):
@@
Step 2 (Taking $L\to\infty$):
@@
where the second step follows from karatzas2012brownian. Combining (A.5), (A.9) and (A.11), we conclude
@@
Step 3 (Taking $m\to\infty$):
@@
Step 4 (Existence of an allocation process satisfying A.14):
@@
By (A.8) and (A.14), there exists a sequence $\{(m_{j},L_{j})\}_{j=1}^{\infty}$ with $(m_{j},L_{j})\to(\infty,\infty)$, under which [^10] It is possible to choose such a sequence since weak convergence can be metrized using the bounded Lipschitz metric.
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ represent the canonical probability space corresponding to $\{z_{a}(\cdot)\}_{a},U$. Specifically, we set $\Omega\equiv C^{d}[0,1]\times C^{d}[0,1]\times[0,1]$, $\mathcal{F}\equiv\mathcal{B}(\Omega)$ and $\mathbb{P}\equiv\mathcal{W}^{d}\otimes\mathcal{W}^{d}\otimes\lambda$, where $C^{d}[0,1]$ represents the space of continuous functions from $[0,1]$ to $\mathbb{R}^{d}$, $\mathcal{W}^{d}$ denotes the $d$-dimensional Wiener measure on $C^{d}[0,1]$, and $\lambda$ is the Lebesgue measure on $[0,1]$. Then, (A.15) implies, by the properties of stable convergence, that [^11] See hausler2015stable for a textbook treatment of stable convergence.
@@
Let $\mathcal{\mathcal{F}}_{\gamma}^{(a)}\subseteq\mathcal{F}$ denote the right-continuous filtration generated by the sample paths of $z_{a}(\cdot)$ between 0 and $\gamma$, and take $\mathcal{G}_{\gamma_{1},\gamma_{0}}\subseteq\mathcal{F}$ to be the augmentation of $\mathcal{\mathcal{F}}_{\gamma_{1}}^{(1)}\lor\mathcal{\mathcal{F}}_{\gamma_{0}}^{(0)}\lor\sigma(U)$ with respect to $(\Omega,\mathcal{F},\mathbb{P})$. In the extended probability space, this gives rise to the extended filtration $\bar{\mathcal{G}}_{\gamma_{1},\gamma_{0}}\equiv\mathcal{G}_{\gamma_{1},\gamma_{0}}\lor\sigma(V)\subseteq\bar{\mathcal{F}}$. Note that the filtration $\bar{\mathcal{G}}_{\gamma_{1},\gamma_{0}}$ inherits the right-continuous and augmented nature of $\mathcal{G}_{\gamma_{1},\gamma_{0}}$. We now argue that $\left\{\bm{Q}(t)\leq\bm{\gamma}\right\}\equiv\{q_{1}(t)\leq\gamma_{1},q_{0}(t)\leq\gamma_{0}\}$ is $\bar{\mathcal{G}}_{\gamma_{1},\gamma_{0}}$ measurable for each $\gamma_{1},\gamma_{0},t\in[0,1]$ such that $\gamma_{1}+\gamma_{0}\geq t$.
@@
Step 5 (Completing the proof):
@@
By the properties of weak convergence (see, e.g., van1996weak), part (iii) of Theorem 1 follows if we show
@@
By karatzas2012brownian, $\lim_{m\to\infty}r_{m}^{(a)}(\epsilon)=0$ for any $\epsilon>0$, so $\lim_{m\to\infty}\bar{r}(m,\epsilon)=2C\epsilon$.
@@ B.1. Proof of Corollary 1
$E\left[\exp\int_{0}^{t}\left\{\sum_{a}\frac{q_{a}(t)}{2}h^{(a)\intercal}I_{a}h^{(a)}\right\}dt\right]\leq\exp\left\{\sum_{a}\frac{1}{2}h^{(a)\intercal}I_{a}h^{(a)}\right\}<\infty.$
@@
Since the processes, $\{x_{a}(\cdot),q_{a}(\cdot)\}_{a}$, are tight, their sample paths lie in a separable metric space $\mathcal{D}$, with an associated Borel sigma-algebra $\mathcal{B}(\mathcal{D})$. This, together with the fact that $V\geq 0$ and $\mathbb{E}[V]=1$, implies, by a version of Le Cam’s third lemma for processes (see, e.g., van1996weak), that
@@ B.2. Proof of Theorem 2
By the definition of weak regularity, the sequence $\sqrt{n}(T_{n}-\bm{\theta}_{0})$ is tight, and consequently, so is the sequence $\sqrt{n}(T_{n}-\bm{\theta}(\bm{h}))=\sqrt{n}(T_{n}-\bm{\theta}_{0})-\bm{h}$. Since the individual elements converge in distribution, it follows that the joint
@@ B.5. Proof of Theorem 4
As in the proof of Theorem 2, $E[V(q_{1}(t),q_{0}(t))]=1$. Furthermore, by the arguments as in that proof again, we also have
@@ Appendix C Auxiliary results for the proof of Theorem 1
are independent. Then, we have that (almost surely): (i) $\varphi_{1,l}=\psi_{1,l}$, and (ii) $\sum_{j\leq l}\varphi_{1,j}\leq\sum_{j\leq l^{\prime}}\varphi_{1,j}$ for all $l^{\prime}\geq l$ with $\sum_{j}\varphi_{1,j}=1$. Property (i) follows from well known properties of regular conditional probabilities (see the proof of Proposition 3 in LeCam1979). Property (ii) follows from (A.3) and the definition of $\varphi_{1,l}$ after noting that $\sum_{j\leq l}\phi_{n,1,j}\leq\sum_{j\leq l^{\prime}}\phi_{n,1,j}$ for all $l^{\prime}\geq l$ and $\sum_{j}\phi_{n,1,j}=1$.
@@
$\mathcal{L}_{k,l}$ $\equiv\sigma\left\{\{s_{1,j}\}_{j},\{s_{0,j}\}_{j},\left\{\tilde{q}_{a,\tilde{k}}^{(m,L)}:\tilde{k}\leq k-1\right\}_{a}\right\},$
$\mathcal{M}_{k,l}$ $\equiv\sigma\left\{\{s_{1,j}\}_{j},\{s_{0,j}\}_{j},\tilde{\bm{\Phi}}^{(k,l)}\right\},$
$\mathcal{H}_{k,l}$ $\equiv\sigma\left\{\{s_{1,j}\}_{j\leq l},\{s_{0,j}\}_{j\leq k\bar{c}-l+1},\tilde{\bm{\Phi}}^{(k,l)}\right\},$
@@
$\tilde{B}_{n,k,l}$ $\equiv\left\{\tilde{q}_{n,1,\tilde{k}}^{(m,L)}\leq\eta_{l},\tilde{q}_{n,0,\tilde{k}}^{(m,L)}\leq\eta_{k\bar{c}-l},\ \forall\ \tilde{k}\leq k-1\right\},$
@@
Also, take $\mathcal{L}_{k,l}^{+}$ and $\mathcal{L}_{k,l}^{-}$ to be the $\sigma$-algebras corresponding to the restriction of $\mathcal{L}_{k,l}$ to $\tilde{B}_{k,l}$ and $\tilde{B}_{k,l}^{c}$, respectively. The quantities $\mathcal{M}_{k,l}^{+},\mathcal{M}_{k,l}^{-}$ and $\mathcal{H}_{k,l}^{+},\mathcal{H}_{k,l}^{-}$ are defined analogously. Observe that $\mathcal{L}_{k,l}^{+}\equiv\mathcal{M}_{k,l}^{+}$. This is because, when $\tilde{B}_{k,l}$ holds, all the $\left\{\tilde{\phi}_{\tilde{k},l}\right\}_{\tilde{k}\leq k-1}$ random variables outside the collection $\tilde{\bm{\Phi}}^{(k,l)}$ necessarily take on the value 0, so they do not provide any additional information.
Define $\vartheta_{k,l}:=\mathbb{E}\left[\tilde{\phi}_{k,l}|\mathcal{L}_{k,l}\right]$ and observe that
$\vartheta_{k,l}$ $=\mathbb{I}\{\tilde{B}_{k,l}\}\mathbb{E}\left[\tilde{\phi}_{k,l}|\mathcal{L}_{k,l}^{+}\right]+\mathbb{I}\{\tilde{B}_{k,l}^{c}\}\mathbb{E}\left[\tilde{\phi}_{k,l}|\mathcal{L}_{k,l}^{-}\right]\ \textrm{a.s.}$
$=\mathbb{I}\{\tilde{B}_{k,l}\}\mathbb{E}\left[\tilde{\phi}_{k,l}|\mathcal{M}_{k,l}^{+}\right]\ \textrm{a.s.}$
The second step uses the fact $\tilde{q}_{a,k}^{(m,L)}$ is almost surely non-decreasing in $k$ as it is the weak limit of $q_{n,a,k}^{(m,L)}$, which is non-decreasing in $k$. Hence, conditional on $\tilde{B}_{k,l}^{c}$, we have $\tilde{\phi}_{k,l}=0\ \textrm{a.s}$, implying $\mathbb{E}\left[\tilde{\phi}_{k,l}|\mathcal{L}_{k,l}^{-}\right]=0\ \textrm{a.s}$.
@@
$\{s_{1,j}\}_{j\leq l},\{s_{0,j}\}_{j\leq k\bar{c}-l+1},\tilde{\bm{\Phi}}^{(k,l)},$
the random variables generating $\mathcal{H}_{k,l}$. Denote this function by $\varphi_{k,l}(\cdot)$.
@@
From the definition of $\phi_{k,l}$ (C.2) and the induction hypothesis, some straightforward algebra reveals that the above random variables are all, in turn,
@@
$\mathbb{P}\left(\left.q_{1,k}^{(m,L)}=\eta_{l}\right|\{s_{a,j}\}_{a,j},\left\{\tilde{q}_{a,k^{\prime}}^{(m,L)}:k^{\prime}\leq k-1\right\}_{a},B_{k,l}\right)$
@@
$=\varphi_{k,l},$
which is the same (a.s.) as $\psi_{k,l}$, the conditional probability of $\left\{\tilde{q}_{1,k}^{(m,L)}=\eta_{l}\right\}$ given $\mathcal{L}_{k,l}^{+}$. At the same time, $\left\{q_{1,k}^{(m,L)}=\eta_{l}\right\}$ never occurs when $B_{k,l}^{c}$ occurs, which matches the fact that the conditional probability of $\left\{\tilde{q}_{1,k}^{(m,L)}=\eta_{l}\right\}$ given $\tilde{B}_{k,l}^{c}$ is 0. So, overall, we conclude that the conditional law of $q_{1,k}^{(m,L)}$ given $\{s_{a,l}\}_{a,l}$ and $\left\{\tilde{q}_{a,k^{\prime}}^{(m,L)}:k^{\prime}\leq k\right\}_{a}$ is the same as the conditional law of $\tilde{q}_{a,k}^{(m,L)}$ given $\mathcal{L}_{k,l}$. Combined with the induction hypothesis, it then follows that the joint law of $\left(\{s_{a,l}\}_{a,l},q_{1,1}^{(m,L)},\dots,q_{1,k}^{(m,L)}\right)$ is equivalent to that of $\left(\{s_{a,l}\}_{a,l},\tilde{q}_{1,1}^{(m,L)},\dots,\tilde{q}_{1,k}^{(m,L)}\right)$. ∎
@@ Appendix D Equivalence of out-of-sample regret
Clearly, $\bm{\delta}_{n}$ must be $\mathcal{I}_{n,1}:=\mathcal{G}_{n,q_{n,1}(1),q_{n,0}(1)}$ measurable.
@@
Suppose Assumptions 1, 2 and 4 hold. Let $\{x_{a}(\cdot),q_{a}(\cdot)\}_{a}$ denote the weak limit of $\{x_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$ under a sequence of policy rules $\{\pi_{n,j}\}_{j}$ in the actual experiment, and $\{\bm{\delta}_{n}\}_{n}$ a sequence of treatment decisions with corresponding out-of-sample regret $W_{n}^{o}(\bm{h})$. Then, there exists a subsequence $\{\bm{\delta}_{n_{k}}\}_{k}$ and a treatment decision $\bm{\delta}$ in the limit experiment, depending only on $\{x_{a}(1),q_{a}(1)\}_{a}$, with out-of-sample regret $W^{o}(\bm{h})$, such that $\lim_{k\to\infty}W_{n_{k}}^{o}(\bm{h})=W^{o}(\bm{h})$ for each $\bm{h}$.
@@ Appendix E Additional results on anytime-valid inference
(ii) For each $a$, $\sqrt{n}\mathbb{E}_{n,\bm{h}}\left[\psi(Y_{i}^{(a)})\right]=I_{a}h+\delta_{n}|h^{(a)}|$, where $\delta_{n}\to 0$ is independent of $\bm{h}$.

### v2

@@ A.2. Proof of Theorem 1
An informal outline of the proof is as follows: We employ dyadic approximations to discretize the empirical allocation processes $\{q_{n,a}(\cdot)\}_{a}$ across both their support (i.e., time) and range. The resulting discrete processes are shown to converge in distribution under $\mathbb{P}_{n,0}$ to a limit (discrete) process that is a function of $(z_{1}(\cdot),z_{0}(\cdot),U)$, where $z_{1}(\cdot),z_{0}(\cdot)$ are independent $d$-dimensional Wiener processes, and $U\sim\textrm{Uniform}[0,1]$ represents exogenous randomization.
Next, we demonstrate that as the discretization becomes arbitrarily fine, the intermediate limit processes converge to a continuous-time allocation process satisfying the criteria in Definition 1. The theory of stable convergence (Häusler and Luschgy, 2015) plays a key role here in ensuring that the informational/measurability constraints satisfied by the empirical allocation processes are preserved during the transition to the continuous-time (and range) limit.
@@
Step 1: Convergence under dyadic approximations.
@@
Step 2: Taking $L\to\infty$.
@@
where the second step follows from Karatzas and Shreve (2012, Lemma 2.4.19). Combining (A.5), (A.9) and (A.11), we conclude
@@
Step 3: Taking $m\to\infty$.
@@
Step 4: Existence of an allocation process satisfying (A.14).
@@
By (A.8) and (A.14), there exists a sequence $\{(m_{j},L_{j})\}_{j=1}^{\infty}$ with $(m_{j},L_{j})\to(\infty,\infty)$, under which [^11] It is possible to choose such a sequence since weak convergence can be metrized, e.g., using the bounded Lipschitz metric.
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ represent the canonical probability space corresponding to $\{z_{a}(\cdot)\}_{a},U$. Specifically, we set $\Omega\equiv C^{d}[0,1]\times C^{d}[0,1]\times[0,1]$, $\mathcal{F}\equiv\mathcal{B}(\Omega)$ and $\mathbb{P}=\mathcal{W}^{d}\otimes\mathcal{W}^{d}\otimes\lambda$, where $C^{d}[0,1]$ represents the space of continuous functions from $[0,1]$ to $\mathbb{R}^{d}$, $\mathcal{W}^{d}$ denotes the $d$-dimensional Wiener measure on $C^{d}[0,1]$, and $\lambda$ is the Lebesgue measure on $[0,1]$. Then, (A.15) implies, by the properties of stable convergence, that [^12] See Häusler and Luschgy (2015) for a textbook treatment of stable convergence.
@@
Let $\mathcal{\mathcal{F}}_{\gamma}^{(a)}\subseteq\mathcal{F}$ denote the right-continuous filtration generated by the sample paths of $z_{a}(\cdot)$ between 0 and $\gamma$, and take $\mathcal{G}_{\gamma_{1},\gamma_{0}}\subseteq\mathcal{F}$ to be the augmentation of $\mathcal{\mathcal{F}}_{\gamma_{1}}^{(1)}\lor\mathcal{\mathcal{F}}_{\gamma_{0}}^{(0)}\lor\sigma(U)$ with respect to $(\Omega,\mathcal{F},\mathbb{P})$. In the extended probability space, this gives rise to the extended filtration $\bar{\mathcal{G}}_{\gamma_{1},\gamma_{0}}\equiv\mathcal{G}_{\gamma_{1},\gamma_{0}}\lor\sigma(V)\subseteq\bar{\mathcal{F}}$. Note that the filtration $\bar{\mathcal{G}}_{\gamma_{1},\gamma_{0}}$ inherits the right-continuous and augmented nature of $\mathcal{G}_{\gamma_{1},\gamma_{0}}$. We now argue that $\left\{\bm{Q}(t)\leq\bm{\gamma}\right\}:=\{q_{1}(t)\leq\gamma_{1},q_{0}(t)\leq\gamma_{0}\}$ is $\bar{\mathcal{G}}_{\gamma_{1},\gamma_{0}}$ measurable for each $\gamma_{1},\gamma_{0},t\in[0,1]$ such that $\gamma_{1}+\gamma_{0}\geq t$.
@@
Step 5: Completing the proof.
@@
By the properties of weak convergence (see, e.g., van der Vaart and Wellner 1996, Chapter 1.12), part (iii) of Theorem 1 follows if we show
@@
By Karatzas and Shreve (2012, Lemma 2.4.19), $\lim_{m\to\infty}r_{m}^{(a)}(\epsilon)=0$ for any $\epsilon>0$, so $\lim_{m\to\infty}\bar{r}(m,\epsilon)=2C\epsilon$.
@@ B.1. Proof of Corollary 1
$E\left[\exp\int_{0}^{1}\left\{\sum_{a}\frac{q_{a}(t)}{2}h^{(a)\intercal}I_{a}h^{(a)}\right\}dt\right]\leq\exp\left\{\sum_{a}\frac{1}{2}h^{(a)\intercal}I_{a}h^{(a)}\right\}<\infty.$
@@
Since the processes, $\{x_{a}(\cdot),q_{a}(\cdot)\}_{a}$, are tight, their sample paths lie in a separable metric space $\mathcal{D}$, with an associated Borel sigma-algebra $\mathcal{B}(\mathcal{D})$. This, together with the fact that $V\geq 0$ and $\mathbb{E}[V]=1$, implies, by a version of Le Cam’s third lemma for processes (see, e.g., van der Vaart and Wellner 1996, Theorem 3.10.7), that
@@ B.2. Proof of Theorem 2
By the definition, the sequence $\sqrt{n}(T_{n}-\bm{\theta}_{0})$ is tight, and consequently, so is the sequence $\sqrt{n}(T_{n}-\bm{\theta}(\bm{h}))=\sqrt{n}(T_{n}-\bm{\theta}_{0})-\bm{h}$. Since the individual elements converge in distribution, it follows that the joint
@@ B.5. Proof of Theorem 4
As in the proof of Theorem 2, $E[V(q_{1}(t),q_{0}(t))]=1$. Furthermore, by the same arguments as in that proof, we also have
@@ Appendix C Auxiliary results for the proof of Theorem 1
are independent. Then, we have that (almost surely): (i) $\varphi_{1,l}=\psi_{1,l}$, and (ii) $\sum_{j\leq l}\varphi_{1,j}\leq\sum_{j\leq l^{\prime}}\varphi_{1,j}$ for all $l^{\prime}\geq l$ with $\sum_{j}\varphi_{1,j}=1$. Property (i) follows from well known properties of regular conditional probabilities (see the proof of Proposition 3 in Le Cam, 1979). Property (ii) follows from (A.3) and the definition of $\varphi_{1,l}$ after noting that $\sum_{j\leq l}\phi_{n,1,j}\leq\sum_{j\leq l^{\prime}}\phi_{n,1,j}$ for all $l^{\prime}\geq l$ and $\sum_{j}\phi_{n,1,j}=1$.
@@
[…]
@@
$B_{n,k,l}$ $\equiv\left\{q_{n,1,\tilde{k}}^{(m,L)}\leq\eta_{l},q_{n,0,\tilde{k}}^{(m,L)}\leq\eta_{k\bar{c}-l},\ \forall\ \tilde{k}\leq k-1\right\},$
@@
[…]
$=\mathbb{I}\{\tilde{B}_{k,l}\}\psi_{k,l}\ \textrm{a.s.}$
[…]
@@
$\{s_{1,j}\}_{j\leq l},\{s_{0,j}\}_{j\leq k\bar{c}-l+1},\tilde{\bm{\Phi}}^{(k,l)},\mathbb{I}\{\tilde{B}_{k,l}\}$
[…]
@@
From the definition of $\phi_{k,l}$ in (C.2) and the induction hypothesis, some straightforward algebra reveals that the above random variables are all, in turn,
@@
$\mathbb{P}\left(\left.q_{1,k}^{(m,L)}=\eta_{l}\right|\{s_{a,j}\}_{a,j},\left\{q_{a,k^{\prime}}^{(m,L)}:k^{\prime}\leq k-1\right\}_{a},B_{k,l}\right)$
@@
$=\varphi_{k,l}\left(\{s_{1,j}\}_{j\leq l},\{s_{0,j}\}_{j\leq k\bar{c}-l+1},\bm{\Phi}^{(k,l)},B_{k,l}\right),$
[…]
@@ Appendix D Equivalence of out-of-sample regret
Note that $\bm{\delta}_{n}$ must be $\mathcal{I}_{n,1}:=\mathcal{G}_{n,q_{n,1}(1),q_{n,0}(1)}$ measurable.
@@
Suppose Assumptions 1, 2 and 4 hold. Let $\{x_{a}(\cdot),q_{a}(\cdot)\}_{a}$ denote the weak limit of $\{x_{n,a}(\cdot),q_{n,a}(\cdot)\}_{a}$ under a sequence of policy rules $\{\pi_{n,j}\}_{j}$ in the actual experiment, and let $\{\bm{\delta}_{n}\}_{n}$ be a sequence of treatment decisions with associated out-of-sample regret $W_{n}^{o}(\bm{h})$. Then, there exists a subsequence $\{n_{k}\}_{k}$ and a limiting treatment decision $\bm{\delta}$ such that:
(i) $\bm{\delta}$ depends only on the terminal states $\{x_{a}(1),q_{a}(1)\}_{a}$; and
(ii) For each $\bm{h}$, $\lim_{k\to\infty}W_{n_{k}}^{o}(\bm{h})=W^{o}(\bm{h})$, where $W^{o}(\bm{h})$ is the out-of-sample regret associated with $\bm{\delta}$ in the limit experiment.
@@ Appendix E Additional results on anytime-valid inference
(ii) For each $a$, $\sqrt{n}\mathbb{E}_{n,\bm{h}}\left[\psi(Y_{i}^{(a)})\right]=I_{a}h^{(a)}+\delta_{n}|h^{(a)}|$, where $\delta_{n}\to 0$ is independent of $\bm{h}$.